A uniform rod is set up so that it can rotate about a perpendicular axis at one of its ends. The length and mass of the rod are 0.825 m and 1.23 kg. What constant-magnitude force acting at the other end of the rod perpendicularly both to the rod and to the axis will accelerate the rod from rest to an angular speed of 6.77 rad/s in 9.69 s?
Answer #1

SOLUTION :


 ⍺ = angular acceleration = change of angular speed / time

= (6.77 - 0)/9.69

= 0.69866 rad/s^2 


Let the applied force be F Newtons at one end of the rod. 


So, moment about the rotational axis = F * L   

= F * 0.825 N-m 


Now, moment about axis = I⍺

I w.r.t. to the axis of rotation at one end 

= m L^2 / 3 

= 1.23 * (0.825)^2 / 3

= 0.27906 kg-m^2


So, 


F* 0.825 = I⍺ = 0.27986 * 0.69866 = 0.19553 kg-m^2/s^2

=> F = 0.19553/0.825 = 0.237 kg-m/s^2 = 0.237 N 

=> F =  0.237 N  (ANSWER)

answered by: Tulsiram Garg
Answer #2

given are

L=0.825m

M=1.23kg

W=6.77rad/sec

T=9.69sec

Use the rotational motion equation linearly

F = m*a

T = I*alpha where I is the moment of inertia of the rod and alpha is the angular acceleration in rad/s^2
I = m*L^2/3

= 1.23*0.825^2/3

= 0.279
Use the rotational equivalent of the linear equation

V = Vi + a*t
ω(t) = ωi + alpha*t

w(t)= 6.77 = 0 + alpha*9.69

alpha = 6.77/9.69

= 0.698 rad/s^2
T = I*alpha

= 0.279*0.698

= 0.194Nm

= 0.825*F

F = 0.194/0.825

F= 0.236N